Solve the following I.V.P. by method of Laplace transform:


(i) 𝑦 β€²β€² + 𝑦 = 𝑓(𝑑), 𝑦(0) = 0, 𝑦 β€² (0) = 0 𝑓(𝑑) = { 2 0 ≀ 𝑑 ≀ 3 3𝑑 βˆ’ 7 3 < 𝑑 < ∞


(ii) 𝑦 β€²β€² + 𝑦 = 𝑓(𝑑), 𝑦(0) = 0, 𝑦 β€² (0) = 0 𝑓(𝑑) = { 𝑑^2 0 ≀ 𝑑 ≀ 1 0 1 < 𝑑 < ∞


1
Expert's answer
2021-10-20T18:17:43-0400

(i)

"y''+y=f(t),y(0)=0,y'(0)=0"

"f(t)=\begin{cases}
 2 &\text{} 0\le t\le 3 \\
 3t-7 &\text{} 3<t<\infty
\end{cases}"

Here f(t) is piecewise continuous on"(0,\infty)", then the derivative theorem exists.

So apply both sides Laplace transformation

"L(y''+y)=L(f(t))"

"L(y''(t))+L(y(t))=L(f(t))"

"s^{2}Y(s)-sY(0)-y'(0)+Y(s)=\int_0^{\infty}{e^{-st}f(t)dt}"

"s^{2}Y(s)-s*0-0+Y(s)=\int_0^{\infty}{e^{-st}f(t)dt}"

"Y(s)(s^{2}+1)=\int_0^{\infty}{e^{-st}f(t)dt}"

"=\int_0^{3}{e^{-st}2dt}+\int_3^{\infty}{e^{-st}(3t-7)dt}"

"=2[\frac{e^{-st}}{-s}]_0^{3}+\int_3^{\infty}{3e^{-st}tdt}-7\int_3^{\infty}{e^{-st}dt}"

"=\frac{-2}{s}[e^{-3s}-1]+3\int_3^{\infty}{te^{-st}dt}-7\int_3^{\infty}{e^{-st}dt}"

"=\frac{-2}{s}[e^{-3s}-1]+3[\frac{te^{-st}}{-s}-\frac{1}{s^{2}}e^{-st}]_3^{\infty}+\frac{7}{s}[e^{-st}]_3^{\infty}"

"=\frac{-2}{s}e^{-3s}+\frac{2}{s}+3[0-0+\frac{e^{-3s}}{s}+\frac{1}{s^{2}}e^{-3s}]+\frac{7}{s}[0-e^{-3s}]"

"=\frac{-2}{s}e^{-3s}+\frac{2}{s}+\frac{3e^{-3s}}{s}+\frac{3e^{-3s}}{s^{2}}+\frac{7e^{-3s}}{s}"

"=e^{-3s}[\frac{-2}{s}+\frac{3}{s}+\frac{3}{s^{2}}-\frac{7}{s}]+\frac{2}{s}"

"=e^{-3s}[\frac{-2s+3s+3-7s}{s^{2}}]+\frac{2}{s}"

"=e^{-3s}[\frac{3-6s}{s^{2}}]+\frac{2}{s}"

"Y(s)=\frac{1}{s^{2}+1}[(e^{-3s}(\frac{3-6s}{s^{2}}))+\frac{2}{s}]"

Where Y(s) is the Laplace transformation of y(t)


(ii)

"y''+y=f(t),y(0)=0,y'(0)=0"

"f(t)=\begin{cases}
 t^{2} &\text{} 0\le t\le 1 \\
 0 &\text{} 1<t<\infty
\end{cases}"

So we apply Laplace transformation on both sides

"L(y''+y)=L(f(t))"

"L(y''(t))+L(y(t))=L(f(t))"

"s^{2}Y(s)-sY(0)-y'(0)+Y(s)=\int_0^{\infty}{e^{-st}f(t)dt}"

"s^{2}Y(s)-0-0+Y(s)=\int_0^{\infty}{e^{-st}f(t)dt}"

Where f(t) is a piecewise continuous function

"Y(s)(s^{2}+1)=\int_0^{1}{t^{2}e^{-st}dt}+\int_1^{\infty}{0*e^{-st}dt}"

"=\int_0^{1}{t^{2}e^{-st}dt}+0"

"=[t^{2}\frac{e^{-st}}{-s}]_0^{1}-\int_0^{1}{2t\frac{e^{-st}}{-s}dt}"

"=[1\frac{e^{-s}}{-s}+0]-\frac{2}{s}[t\frac{e^{-st}}{-s}-\frac{1}{s^{2}}e^{-st}]_0^{1}"

"=-\frac{e^{-s}}{s}-\frac{2}{s}[\frac{e^{-s}}{-s}-\frac{e^{-s}}{s^{2}}+\frac{1}{s^{2}}]"

"=-\frac{e^{-s}}{s}+\frac{2e^{-s}}{s^{2}}+\frac{2e^{-s}}{s^{3}}-\frac{2}{s^{3}}"

"=\frac{e^{-s}}{s}[-1+\frac{2}{s}+\frac{2}{s^{2}}]-\frac{2}{s^{3}}"

"Y(s)=\frac{e^{-s}}{(s^{2}+1)s}[-1+\frac{2}{s}+\frac{2}{s^{2}}]-\frac{2}{s^{3}(s^{2}+1)}"

Where Y(s) is the Laplace transformation of y(t).